OpenStudy (anonymous):

Use a transformation to solve the equation. m + 12.6 = –24 A. m = –36.6 B. m = –12.2 C. m = 12.2 D. m = 36.6

OpenStudy (kaylardz0405):

-24 -12.6

OpenStudy (anonymous):

@IGreen what happened to him?

OpenStudy (studytheworld):

m=-36.6

OpenStudy (anonymous):

No direct answers. m + 12.6 = –24 Subtract 12.6 m = -24 - 12.6 m= -24 + -12.6 m= -36.6